Please help me with this, I need as much help as I can get
If f(x)=2^x,describe the transformation of the curve for g(x)=-2^x-2

Answer by stanbon(48535)

If f(x)=2^x,describe the transformation of the curve for g(x)=-2^x-2
------------------
Looking at g(x) and comparing it with f(x):
Start with f(x)=2^x
Then f(x)=2^(x-2) which moves the 1st curve two to the right.
Then f(x)=-2^(x-2) which reflects the last curve in the x axis
and double the y values.
The resulting curve is g(x)
Cheers,
Stan H.